# Optimal Protocols for Continual Learning via Statistical Physics and Control Theory

### v1

@@ A.1 Generalisation error as a function of the order parameters
$\begin{split}\varepsilon_{t}\left(\bm{W},\bm{V},\bm{W}_{*}\right)&=\frac{1}{2\pi}\sum_{k,h}v_{k}^{(t)}v_{h}^{(t)}\arcsin\frac{Q_{kh}}{\sqrt{1+Q_{kk}}\sqrt{1+Q_{hh}}}+\frac{1}{2\pi}\arcsin\frac{S_{tt}}{1+S_{tt}}\\ &\quad-\frac{1}{\pi}\sum_{k}v_{k}^{(t)}\arcsin\frac{M_{kt}}{\sqrt{1+Q_{kk}}\sqrt{1+S_{tt}}}\;.\end{split}$ (21)
@@ A.2 Ordinary differential equations for the forward training dynamics
where we stress the difference between $t_{c}$, the task selected for training at epoch $\mu$, and $t$, the task for which we compute the overlap. We define a “training time” $\alpha=\mu/N$ and take the infinite-dimensional limit $N\rightarrow\infty$. The parameter $\alpha$ becomes continuous and $M_{kt}$ concentrates to the solution of the following ODE:
@@
The initial condition for $\mathbb{Q}(0)$ is chosen to reproduce the random initialisation of the SGD algorithm. In particular, the initial first-layer weights and readout weights are drawn i.i.d. from a normal distribution with variances of $10^{-3}$ and $10^{-2}$, respectively.
@@ A.3 Informal derivation of Pontryagin maximum principle
Let us consider the augmented cost function
@@
Note that we do not consider variations with respect to $\mathbb{Q}(0)$ as this quantity is fixed by the initial condition $\mathbb{Q}(0)=\mathbb{Q}_{0}$. Finally, minimizing the cost function with respect to the control ${\bm{u}}$, we get the optimality condition in Eq. 9 of the main text.
@@ A.4 Optimal control framework
To determine the optimal control, we iterate Eqs. 4, 7, and 9 of the main text until convergence (Bechhoefer, 2021). Let us consider first the case where the control is the current task $t_{c}(\alpha)$, such that $t_{c}(\alpha)=t$ if the network is trained on task $t\in\{1\,,\ldots\,,T\}$ at training time $\alpha$. For simplicity, we focus on the case $T=2$, but the following discussion is easily generalised to any $T$. In particular, since here ${u}(\alpha)=t_{c}(\alpha)$ the evolution equation 4 can be written as
@@
Then, we start from a guess for the control variable $t_{c}(\alpha)$. We integrate Eq. 40 forward, obtaining the trajectory $\mathbb{Q}(\alpha)$ for $\alpha\in(0,\alpha_{F})$. Then, we integrate the backward equation 41, starting from the final condition 42, obtaining the trajectory ^$\hat{\mathbb{Q}}(\alpha)$ for $\alpha\in(0,\alpha_{F})$. Then, the control variable can be updated using Eq. 44 and used in the next iteration of the algorithm. These equations 40, 41, and 44 are iterated until convergence.
We next consider the joint optimisation of the learning rate schedule $\eta(\alpha)$ and the task protocol $t_{c}(\alpha)$. The optimality condition 9 can be written as
@@ Appendix C Supplementary figures
Fig. 4 compares the values of the loss at the end of training, averaged on both tasks, for different task-selection strategies. In particular, it highlights the performance gap between the four replay strategies at constant learning rate considered in the main text (no-replay, interleaved, optimal and pseudo-optimal) and the strategy that simultaneously optimise over task-selection and learning rate.
@@
Figure 9: Adopting an optimal learning rate schedule leads to major perfomance improvement. Average loss on both tasks at the end of the second training phase as a function of task similarity $\gamma$ under the same setting and parameters as Fig. 3 of the main text. The top four lines correspond to different strategies at constant learning rate $\eta=1$: no replay (purple crosses), optimal replay (red dots), interleaved (blue squares), pseudo-optimal replay (cyan dashed line). The bottom curve (brown plus signs) corresponds to jointly optimal replay and learning rate schedules (see Fig. 4).

### v2

@@ A.1 Generalisation error as a function of the order parameters
$\begin{split}\varepsilon_{t}\left(\bm{W},\bm{V},\bm{W}_{*}\right)&=\frac{1}{\pi}\sum_{k,h}v_{k}^{(t)}v_{h}^{(t)}\arcsin\frac{Q_{kh}}{\sqrt{1+Q_{kk}}\sqrt{1+Q_{hh}}}+\frac{1}{\pi}\arcsin\frac{S_{tt}}{1+S_{tt}}\\ &\quad-\frac{2}{\pi}\sum_{k}v_{k}^{(t)}\arcsin\frac{M_{kt}}{\sqrt{1+Q_{kk}}\sqrt{1+S_{tt}}}\;.\end{split}$ (22)
@@ A.2 Ordinary differential equations for the forward training dynamics
where we stress the difference between $t_{c}$, the task selected for training at step $\mu$, and $t$, the task for which we compute the overlap. We define a “training time” $\alpha=\mu/N$ and take the infinite-dimensional limit $N\rightarrow\infty$. The parameter $\alpha$ becomes continuous and $M_{kt}$ concentrates to the solution of the following ODE:
@@
The initial condition for $\mathbb{Q}(0)$ is chosen to reproduce the random initialisation of the SGD algorithm. In particular, the initial first-layer weights and readout weights are drawn i.i.d. from a normal distribution with variances of $10^{-3}$ and $10^{-2}$, respectively. A thorough analysis of the validity of this ODE description is provided in Veiga et al. (2022), where the authors study the crossover between narrow and infinitely wide networks, clarifying the connection with the so-called mean-field or hydrodynamic regime (Mei et al., 2018; Chizat & Bach, 2018; Rotskoff & Vanden-Eijnden, 2022).
@@ A.3 Informal derivation of Pontryagin maximum principle
Let us consider the augmented cost functional
@@
Note that we do not consider variations with respect to $\mathbb{Q}(0)$ as this quantity is fixed by the initial condition $\mathbb{Q}(0)=\mathbb{Q}_{0}$. Finally, minimizing the cost functional with respect to the control ${\bm{u}}$, we get the optimality condition in Eq. 10 of the main text.
@@ A.4 Optimal control framework
To determine the optimal control, we iterate Eqs. 5, 8, and 10 of the main text until convergence (Bechhoefer, 2021). Let us consider first the case where the control is the current task $t_{c}(\alpha)$, such that $t_{c}(\alpha)=t$ if the network is trained on task $t\in\{1\,,\ldots\,,T\}$ at training time $\alpha$. For simplicity, we focus on the case $T=2$, but the following discussion is easily generalised to any $T$. In particular, since here ${u}(\alpha)=t_{c}(\alpha)$ the evolution equation 5 can be written as
@@
Then, we start from a guess for the control variable $t_{c}(\alpha)$. We integrate Eq. 41 forward, obtaining the trajectory $\mathbb{Q}(\alpha)$ for $\alpha\in(0,\alpha_{F})$. Then, we integrate the backward equation 42, starting from the final condition 43, obtaining the trajectory ^$\hat{\mathbb{Q}}(\alpha)$ for $\alpha\in(0,\alpha_{F})$. Then, the control variable can be updated using Eq. 45 and used in the next iteration of the algorithm. These equations 41, 42, and 45 are iterated until convergence.
We next consider the joint optimisation of the learning rate schedule $\eta(\alpha)$ and the task protocol $t_{c}(\alpha)$. The optimality condition 10 can be written as
@@ C.1 Additional results in the synthetic framework
Fig. 9 compares the values of the loss at the end of training, averaged on both tasks, for different task-selection strategies. In particular, it highlights the performance gap between the four replay strategies at constant learning rate considered in the main text (no-replay, interleaved, optimal and pseudo-optimal) and the strategy that simultaneously optimise over task-selection and learning rate. Additionally, we consider exponential and power law learning rate schedules, in combination with interleaved replay protocol. For each value of task similarity $\gamma$, we optimize over the schedule parameters via grid search. We still find a performance gap with respect to the optimal strategy, which highlights the relevance of the joint optimization of training protocols.
Fig. 10 illustrates a continual learning setting with $T=3$ tasks. The student is a two-layer neural networks with $K=3$ hidden units, and a different readout is trained for each task. In the initial training phase, the student is trained on task $1$ up to time $\alpha=1000$, when the loss reaches $\sim 10^{-6}$. In the second phase, the student must learn tasks $2$ and $3$ without forgetting task $1$. Panel a) shows the losses of the optimal strategy as a function of time for the three tasks and their average, during the second training phase. The optimal strategy is represented by a colour bar in the upper panel. It consists in an initial phase where only the new tasks $2$ and $3$ are presented to the network, and a second phase where task $1$ is replayed. Despite its complicated structure, it shares some similarities with the pseudo-optimal strategy described in the main text. Specifically, task 1 is replayed only when its loss is comparable to the losses on the other two tasks. Panel b) shows the average loss as a function of time, comparing different task-selection protocols. In particular, we consider
  • interleaved protocols, where two or all three tasks are alternated during training;
  • sequential protocols, where tasks are presented in distinct blocks without being replayed;
  • a shuffled protocol, that preserves the relative fraction of samples from each tasks obtained from the optimal strategy but presents them in a randomised order.
The final performance of the optimal strategy surpasses all the aforementioned approaches. Notably, as the number of tasks grows, the number of possible heuristic strategies expands significantly, making it difficult to identify effective solutions through intuition alone. This highlights the importance of a theoretical framework for systematically determining the optimal strategy.
@@
Figure 9: Adopting an optimal learning rate schedule leads to major perfomance improvement. Average loss on both tasks at the end of the second training phase as a function of task similarity $\gamma$ under the same setting and parameters as Fig. 3 of the main text. The top four lines correspond to different strategies at constant learning rate $\eta=1$: no replay (purple crosses), optimal (red dots), interleaved (blue squares), pseudo-optimal (cyan dashed line). The bottom three curves correspond to different annealing schemes for the learning rate. Up-facing orange triangles correspond to interleaved replay with power law annealing $\eta(\alpha)=\eta_{0}(1-\alpha/\alpha_{f})^{\beta}$, while down-facing green triangles indicate exponential annealing $\eta(\alpha)=\eta_{0}\exp(-\alpha/\alpha_{0})$. For both annealing schedules, the scalar parameters $\eta_{0}$, $\beta$, and $\alpha_{0}$ are optimized by grid search for each value of $\gamma$. Finally, the brown plus signs correspond to jointly optimal replay and learning rate schedules (see Fig. 4).
Figure 10: Optimal replay schedule for $T=3$ tasks. The student network has $k=3$ hidden units. In the initial phase, $\alpha=[0,1000]$ the network is trained on task 1 until convergence (loss is ${}10^{-6}$). In the second phase $\alpha=[1000,1025]$, we determine the optimal replay strategy. In both phases the learning rate is constant $\eta=1$. The tasks are chosen such that the overlaps are $S_{1,2}=0.5$, $S_{2,3}=0.5$, and $S_{1,3}=0$. Panel a) shows the optimal task protocol and the evolution of the loss over the three tasks. The result is a complicated replay strategy, though it shares some similarities with the pseudo-optimal strategy described in the main text. Specifically, task 1 is replayed only when its loss is comparable to the losses on the other two tasks. Panel b) compares the optimal strategy with different heuristics, all at $\eta=1$. “Interleaved $(1-2-3)$” is an interleaved strategy containing all tasks in equal proportion. “Interleaved $(2-3)$” is an interleaved strategy containing tasks $(2-3)$ in equal proportion. “Shuffled” contains the same task proportions as in the optimal strategy, but in random order (showing that the structure of the optimal replay strategy matters). “Sequential $(2-3)$” corresponds to the sequential strategy with $t=2$ in the first half and $t=3$ in the second. Similarly for “Sequential $(3-2)$”. “Sequential $(3-2-1)$” has $t=3$ in the first third of the replay sequence, then $t=2$ in the second third and $t=1$ in the last third.
C.2 Additional results on real dataset
We run additional experiments on real data using the simulation setup detailed in Sec. 3.2. We consider the CIFAR10 dataset Krizhevsky et al. (2009) and create two classification tasks taking as task 1 the classes with odd labels and the others as task 2. Fig. 11 shows the results of the training curve for the two tasks according to the three strategies and the final figure compares the average loss throughout the training steps. Notice that the observation reported in the controlled Fashion MNIST experiment are still valid in this scenario.
Figure 11: Continual learning on CIFAR10. Training curves of No-Replay, Interleaved, and Pseudo-Optimal learning strategies (from left to right) on CIFAR10. The two tasks are obtained partitioning the dataset according to the parity of the labels and training using online stochastic gradient descent. The final panel (rightmost) shows the average loss for the three strategies. The inset zooms on the final stage of learning.
As already observed in the main text, the performance of the Pseudo-Optimal learning strategies appears to interpolate between No-Replay and Interleaved. In Fig. 12 we highlight the differences between Pseudo-Optimal and the alternative strategies. We see a up to 3% improvement in performance during learning. Contrarily to the main text, we observe also a region where Pseudo-Optimal appears sub-performing with respect to the Interleaved strategies, however this is limited to a less than 0.5% decrease in loss.
Figure 12: Learning difference on CIFAR10. This figure complement Fig. 11, highlighting the differences between the strategies.